OpenStudy (anonymous):

Can someone Help me with this pls! H(x)=ln.sqrt(a^2 -z^2/a^2 +z^2)

OpenStudy (anonymous):

what the question is?

OpenStudy (anonymous):

Differentiate the function!

OpenStudy (anonymous):

partial derivative?

OpenStudy (anonymous):

I guess Inverse functions

OpenStudy (anonymous):

which class are you in? cal1,2,3?

OpenStudy (anonymous):

cal-2

OpenStudy (anonymous):

ok, to me, I break the ln and sqr into 2 parts. I have H(x) = ln (sqr(a^2 -z^2) ) - ln (sqr(a^2 + z^2) ). then take derivative each of them. don't forget each is 3 layers function

OpenStudy (anonymous):

at cal2, you just have 1 variable, I assume that your a is constant, so, just take derivative respect to z

OpenStudy (anonymous):

HOW CAN YOU BREAK THE sqrt into 2?

OpenStudy (anonymous):

I mean, is it possible?

OpenStudy (anonymous):

\[\sqrt{\frac{ a^2-z^2 }{ a^2+z^2 }}= \frac{ \sqrt{a^2-z^2} }{ \sqrt{a^2+z^2} }\]

OpenStudy (anonymous):

is it ok?

OpenStudy (anonymous):

Thanks man :) btw what do U mean by 3 layers function?

OpenStudy (anonymous):

you have ln-sqrroot- sqr . that is 3 layers

OpenStudy (anonymous):

when take derivative, go from outside to inside, i mean from left to right.

OpenStudy (anonymous):

unfan me please, I just stop by for fun,not staying here long, try to become a gentlement

OpenStudy (anonymous):

OK! but I still didn't get the case"3 layers functions"! becoz we have ln & sqrt there so it must be 2 layers!!!

OpenStudy (anonymous):

hey, how about z^2?

OpenStudy (anonymous):

when you "touch" it, you must take (z^2)' to get ... something, I don't know

OpenStudy (anonymous):

it's inside the root , so the whole function under the root must be one layer itself! am I wrong?

OpenStudy (anonymous):

unfortunately, you are not right. if z is just z, you are right, but z is z^2 , you must consider it a function of z

OpenStudy (anonymous):

So ln, a^2, z^2 each considered as a function here?!! btw thx in advance ;)

OpenStudy (anonymous):

I mean as a layer!

OpenStudy (anonymous):

1 personal question: are you taking cal-3? which university?

OpenStudy (anonymous):

I do the first part for you [ln(sqr)(a^2 -z^2 ]= [1/sqr (a^2-z^2) ]*[ (sqr(a^2 -z^2)]' . the second part of this part ask you take derivative of sqr, right? I break it again to count just that part. you must time to the first part then. [sqr(a^2-z^2)]' = 1/2sqr(a^2-z^2) * (a^2-z^2)' . and (a^2-z^2)' = 2z. is it right?

OpenStudy (anonymous):

yeap. not university, just college. community college of Philadelphia

OpenStudy (anonymous):

well, thank you so much for your help :) by the way you look one of those timid girls who are running away from guys, aren't you? (bcoz U asked me to unfollow you!)

OpenStudy (anonymous):

nope, I am a gentlemen. I do it because there are so many people become my fan and waiting for help when they need help while I don't have time for them. I don't want you become one of them, ask for help and then wait hopelessly.

OpenStudy (anonymous):

I have to go now, if I see you when you need help AND I have time, definitely I am there. bye bye

OpenStudy (anonymous):

yeah, I see! have a good time (it is night here but it must be afternoon there...! i usually come for questions at nights) ;) bye now